#1) Please complete the problems showing all calculations and equations on which you base your answers. Use -9.8 m/s2 as the value for the acceleration due to gravity. Please be advised below questions is for Biomechanics class. THANK YOU IN ADVANCE!

#1) Please complete the problems showing all calcu
Answer #1

apply momentum conservatiion in both direction

P i = P f ..................................in x dire. s o

take right as positive so

20*9.7 *sin( 48) + 21*7.5 *sin ( 31) - 18*5.2* cos( 20) = ( 20+21+18) Vx

Vx = 2.33 m/s in +x

apply same in Y tkae up as positive so

-20*9.7 *cos( 48) + 21*7.5 *cos( 31) +18*5.2* sin( 20) = ( 20+21+18) Vy

Vy = 0.631 m/s up

so

magnitude of final speed

V = sqrt( 2.33^2 + 0.631^2) = 2.4139 m/s answer

direction = atan( 0.631 / 2.33)

= 15.15 of norht( up) of east ( right)

part2

F = \mu* M*g   

a = \mu*g = 0.07* 9.8 = 0.686 m/s^2

v^2 = u^2 +2*a*s

0 = 2.4139 ^2 - 2* 0.686*s

s = 4.25 m ...................answer